Hello,
I need help finding the length of a hypotenuse of an isosceles right triangle if each leg is 33cm long.
Thank you

Answer by checkley75(3666)

A RIGHT TRIANGLE WITH SIDES OF 33CM HAS AN HYPOTENUSE OF:
A^2+B^2=C^2 WHERE C IS THE HYPOTENUSE
H^2=33^3+33^2
H^2=1089+1089
H^2=2178
H=46.67 CM. IS THE HYPOTENUSE OF THIS TRIANGLE.